# Benchmarking and Improving Bird’s Eye View Perception Robustness in Autonomous Driving

### v1

@@ 2.2 LiDAR-Based 3D Perception
LiDAR, with its precision in capturing spatial relationships using laser beams, has paved the way for breakthroughs in 3D perception, central to applications like autonomous driving [11]. Two primary tasks have gained prominence: 3D object detection and LiDAR semantic segmentation, both of which have inherent connections to BEV perception [30, 5, 33, 23, 13]. In the realm of 3D object detection, the focus has been on optimally representing LiDAR point cloud data [32]. Point-based approaches, such as those presented in [88, 89, 90, 91], shine in preserving the innate geometry of point clouds, capturing local structures and patterns. Meanwhile, voxel-based strategies, like [92, 93, 94], convert the irregular point clouds into structured grids, relying on sparse convolution techniques [92] to handle non-empty voxels efficiently. Pillar-based techniques, highlighted by works like [29, 95], offer a trade-off between detection accuracy and computational speed by fine-tuning the vertical resolution. Additionally, hybrid approaches, such as [96, 97], merge the strengths of both point and voxel representations to derive more enriched features. On the other hand, semantic segmentation techniques often pivot on the representation choice. Raw point methods, like [98, 35], emphasize the direct usage of irregular point clouds, while range view approaches, showcased in [99, 100, 36, 108, 102], convert these point clouds into 2D grids. This conversion aligns closely with BEV perception, transforming 3D data into a top-down perspective, essential for many applications. Further refining this idea are bird’s eye view techniques, exemplified by [103], which offer a direct 2D top-down representation. Voxel-centric methods, such as [104], maintain the 3D spatial structure, often outperforming other singular modalities. Modern research, like [105, 107, 106, 34, 37, 38, 39, 18], pushes the boundaries by exploring the fusion of multiple views, seeking to harness the complementary strengths of different representations. In essence, while LiDAR-based 3D perception methodologies, especially those linked with BEV perception, have exhibited significant promise, their resilience in real-world conditions warrants deeper exploration and validation.
@@ 3.3 Camera-LiDAR Fusion
The BEV paradigm streamlines the fusion of features from a variety of input modalities. While some algorithms focus on crafting BEV representations solely from images, a notable fraction of the literature, including works like [66, 67, 81, 69, 82, 83], advocates for a unified BEV space. This aligns features extracted from both images and point clouds. We delve deep into the performance of such multi-modal fusion algorithms, especially under circumstances where images are corrupted, yet the LiDAR mechanism remains benign. Furthermore, we address a common scenario where the model is trained using multi-modal inputs but deployed on vehicles equipped with only one of the sensors. To assess robustness, we evaluate the model’s performance under conditions of complete sensor failure, where either the camera or LiDAR is missing.
@@ 3.4 BEV View Transformation
The body of works in BEV transformation can be divided based on the use of depth estimation techniques [6]. One faction, as discussed in [3, 22, 51, 65], embeds a distinct depth-estimation branch within their systems. Given the inherent challenges in predicting 3D bounding boxes from images, these models first forecast a per-pixel depth map. This map then serves as a compass, guiding image features to their rightful 3D coordinates. The subsequent BEV transformation process often follows a bottom-up approach, as depicted in [4]. On the other side of the spectrum are models leveraging pre-defined object queries [2, 4] or lean proposals [55, 54] to collate 2D features in a top-down manner. While both these paradigms have demonstrated their prowess on benign datasets, we further expand the horizon by examining their efficacy on corrupted data.
@@ 4.1 Dataset Generation
We also consider scenarios simulating complete sensor blackouts in our camera-LiDAR fusion algorithms. Every pixel is set to zero when simulating the camera’s absence. To emulate the missing of LiDAR readings, only the data points within a [[-[ -45, 45]]]] degree frontal field-of-view (FOV) are retained. Such a design choice is rooted in our observations that multi-modal trained models suffer from significant performance drops when LiDAR readings are entirely absent.
@@ 5.2 Camera-Only Benchmark
Our investigations further highlight a potential vulnerability in resilience rates across various corruptions. Even though the mCE metric displays a linear correlation between the nuScenes and nuScenes-C datasets, the mRR metric elucidates notable disparities among models with comparable baseline performance. This suggests potential overfitting of some models to the nuScenes dataset, thereby compromising their adaptability to the nuScenes-C dataset. For instance, despite Sparse4D [55] outperforming DETR3D [4] on the “clean” dataset, it falls short in terms of mRR metrics across all corruption categories. Moreover, DETR3D performs exceptionally well in Dark conditions, in stark contrast to BEVerse (swin-t). Despite its strong performance in clean conditions, BEVerse (swin-t) achieves only 12% relative performance in the dark. Therefore, a comprehensive assessment of cutting-edge models is essential for a complete evaluation of their capabilities.
As shown in Table II, to gain deeper insights into the model’s robustness, we dissected BEV models based on components like training strategies (e.g., pre-training and CBGS [46] resampling), model architectures (e.g., backbone), and learning techniques (e.g., temporal cue learning).
@@
TABLE VI: Benchmark results for robustness under sensor failures. The models are trained using multi-modal inputs while tested using single sensor inputs.
@@ 5.3.1 Camera Sensor Corruptions
We studied scenarios where cameras are impaired while LiDAR operates optimally, a frequent occurrence in real-world conditions. For instance, LiDAR point cloud capture remains largely unhampered by lighting variations, whereas camera captures can degrade under limited light. Intentionally, we excluded conditions like Snow and Fog, as they could introduce noise to both camera and LiDAR readings. Results of these studies are depicted in Table V. Interestingly, multi-modal fusion models maintain high performance even when the camera data is compromised. When provided with benign LiDAR and degraded camera inputs, BEVFusion [66] consistently outperforms its LiDAR-only counterpart, with a notably higher NDS score of 0.6928, across most types of camera corruptions, except Dark. This affirms the efficacy of using LiDAR data even when the camera data is suboptimal.
@@ 5.4 Validity Assessment
Since the corruption images are synthesized digitally, it is important to study how close they are compared to real-world corruption. To study the validity of synthesized images, we conducted two experiments, including the pixel distribution study and corruption-augmented training.
@@ 5.4.1 Pixel Distributions
Assuming that a corruption simulation is realistic enough to reflect real-world situations, the distribution of a corrupted “clean” set should be similar to that of the real-world corruption set. We validate this using ACDC [110], nuScenes [1], Cityscapes [112], and Foggy-Cityscapes [111], since these datasets contain real-world corruption data and clean data collected by the same sensor types from the same physical locations. We simulate corruptions using “clean” images and compare the distribution patterns with their corresponding real-world corrupted data. We do this to ensure that there is no extra distribution shift from aspects like sensor difference (e.g., FOVs and resolutions) and location discrepancy (e.g., environmental and semantic changes). As illustrated here, the pixel distributions of our synthesized images exhibit a high degree of resemblance to those of real-world data, thereby affirming the dataset’s validity from a pixel statistical perspective.
@@ 5.4.2 Corruption-Augmented Training
Furthermore, if the corruption simulation is realistic enough to reflect real-world situations, a corruption-augmented model should achieve better generalizability than the “clean” model when tested on real-world corruption datasets. Also, the corruption-augmented model should also show better performance on the clean dataset. We validate this using nuScenes, nuScenes-Night, and nuScenes-Rain. We adopt FCOS3D [9] as the baseline and train the model with corruption augmentation. For nuScenes-Night and nuScenes-Rain, we train the model on the Day-train and Dry-train split and evaluate on the Day-val, Night-val, Dry-val, and Rain-val split. The results can be seen in Table VIII. We observe that using synthesized images as the data augmentation strategy successfully improves the cross-domain robustness. Specifically, in day-to-night domain transitions, we observe a significant performance drop from 0.3867 to 0.0854 in the baseline model due to the large domain gap. However, when trained with corruption augmentation, the model’s cross-domain performance improves by 45.8%, thereby validating the validity of our synthesized images.
@@ 5.5 Robustness Improvement
Build-upon our experiment, we further study how to fundamentally improve the robustness of BEV perception models. In Subsection 5.5.1, we simply use the corruption as data augmentation as in Subsection 5.4.2. In Subsection 5.5.2, we resort to advanced CLIP [113] as the model backbone since CLIP is well-known to show promising out-of-distribution generality [116, 115].
@@
5.5.1 Corruption-Augmentation Training
We investigate the utilization of corruption as a data augmentation strategy during the training phase. To systematically assess the effectiveness of this approach, we apply corruption augmentation to five models within our benchmark, with results detailed in Table IX. Our findings indicate that corruption augmentation substantially enhances performance against semantic corruptions within our proposed dataset, particularly for corruption types that originally posed significant challenges to the models (e.g., Motion Blur, Snow). However, augmenting training data with sensor corruption scenarios, such as the omission of camera information, does not significantly improve robustness. This outcome underscores the necessity for future research to develop more sophisticated modules capable of handling incomplete input scenarios.
@@ 5.5.2 Foundation Model Backbone
Recent studies [116, 115] have demonstrated that foundation models, trained on internet-scale datasets in an unsupervised manner, exhibit notable generality compared to models trained on conventional datasets (e.g., ImageNet [117]). Inspired by these findings, we explore the potential to transfer the generality of foundation models to BEV perception tasks. Specifically, we investigate three distinct approaches for utilizing the CLIP backbone in our models: (1) freezing the backbone while training only the detector head; (2) fine-tuning both the backbone and the detector head; (3) freezing the backbone and training the detection head, then fine-tuning the whole model together, since previous work [118] shows that end-to-end fine-tuning, even though improve in-distribution performance, can compromise robustness on out-of-distribution dataset.
The three methods are shown in Figure 4 and the results are shown in Table  X. The first observation is CLIP is not well optimized towards BEV perception tasks, indicated by the low benign performance when we freeze the CLIP backbone. Additionally, end-to-end fine-tuning with randomly initialized detection heads leads to minimal improvement. Interestingly, when applying corruption augmentation, the CLIP backbone shows little improvement to the baseline model. The improvement in mRR is only 0.56 and, when equipped with corruption augmentation, the mRR is even lower. Finally, our two-stage training can effectively improve the performance while also transferring the robustness of CLIP towards BEV perception tasks, especially with equipped with corruption augmentation. The robustness improvement surpasses the end-to-end fine-tuned CLIP by a notable margin. For instance, the NDS score under Dark, Fog, and Snow improve by 23.1%, 11.8%, and 15.8% respectively.
@@ 6.2 Model Pre-Training
- Pre-training improves robustness across a wide range of semantic corruptions while does not help with temporal corruptions. The effectiveness of these strategies for improving model robustness is illustrated in Figure 6a and Figure 6b, where models that utilize pre-training largely outperform those not. For controlled comparison, we re-implement the BEVDet (r101) [3] model using the FCOS3D [9] checkpoint as initialization. Our results, presented in Figure 10, show that pre-training can significantly improve mRR across a wide range of corruptions (except Fog) even if it has lower “clean” NDS (0.3780 vs. 0.3877). Specifically, under Color Quant, Motion Blur, and Dark corruptions, the mRR metric improves by 22.5%, 17.2%, and 27.8%, respectively. It is worth noting that pre-training mainly improves most semantic corruptions and does not improve temporal corruptions. Even though, the pre-trained BEVDet still largely lags behind those depth-free counterparts. Therefore, we can conclude that pre-training together with the depth-free bird’s eye view transformation provides models with strong robustness. More recently, M-BEV [114] propose masked pre-training task to enhance robustness under incomplete sensor input. We compare a masked pre-trained PETR with the corresponding baseline. The results can be seen in Table XI. We find M-BEV can most effectively improve robustness towards incomplete sensor output, and reveal the promise of masked image modeling pre-train in BEV perception tasks.
@@ 6.3 Temporal Fusion
- Temporal fusion has the potential to yield better absolute performance under corruptions. Fusing longer temporal information largely helps with robustness. We are particularly interested in examining how models utilizing temporal information perform under temporal corruptions. We find SOLOFusion [65] which fuses wider and richer temporal information performs extremely well compared to its short-only and long-only versions. In terms of Camera Crash, the short-only and long-only versions have close resilience rate performance (65.04 vs. 65.13). However, the fusion version improves to 70.73, which is the highest among all the candidate models. Similarly, the fusion version improves the resilience rate by almost 10% compared to the other two versions under Frame Lost corruption. Moreover, the RR metric of its long-only version outperforms its short-only counterpart on a wide range of corruption types, which indicates the great potential of utilizing longer temporal information.

### v2

@@ 2.2 LiDAR-Based 3D Perception
LiDAR, with its precision in capturing spatial relationships using laser beams, has paved the way for breakthroughs in 3D perception, central to applications like autonomous driving [11]. Two primary tasks have gained prominence: 3D object detection and LiDAR semantic segmentation, both of which have inherent connections to BEV perception [30, 5, 33, 23, 13]. In the realm of 3D object detection, the focus has been on optimally representing LiDAR point cloud data [32]. Point-based approaches, such as those presented in [88, 89, 90, 91], shine in preserving the innate geometry of point clouds, capturing local structures and patterns. Meanwhile, voxel-based strategies, like [92, 93, 94], convert the irregular point clouds into structured grids, relying on sparse convolution techniques [92] to handle non-empty voxels efficiently. Pillar-based techniques, highlighted by works like [29, 95], offer a trade-off between detection accuracy and computational speed by fine-tuning the vertical resolution. Additionally, hybrid approaches, such as [96, 97], merge the strengths of both point and voxel representations to derive more enriched features. On the other hand, semantic segmentation techniques often pivot on the representation choice. Raw point methods, like [98, 35], emphasize the direct usage of irregular point clouds, while range view approaches, showcased in [99, 100, 36, 108, 102], convert these point clouds into 2D grids. This conversion aligns closely with BEV perception, transforming 3D data into a top-down perspective, which is essential for many applications. Further refining this idea are bird’s eye view techniques, exemplified by [103], which offer a direct 2D top-down representation. Voxel-centric methods, such as [104], maintain the 3D spatial structure, often outperforming other singular modalities. Modern research, like [105, 107, 106, 34, 37, 38, 39, 18], pushes the boundaries by exploring the fusion of multiple views, seeking to harness the complementary strengths of different representations. In essence, while LiDAR-based 3D perception methodologies, especially those linked with BEV perception, have exhibited significant promise, their resilience in real-world conditions warrants deeper exploration and validation.
@@ 3.3 Camera-LiDAR Fusion
The BEV paradigm streamlines the fusion of features from a variety of input modalities. While some algorithms focus on crafting BEV representations solely from images, a notable fraction of the literature, including works like [66, 67, 81, 69, 82, 83], advocates for a unified BEV space. This aligns features extracted from both images and point clouds. We delve deep into the performance of such multi-modal fusion algorithms, especially under circumstances where images are corrupted, yet the LiDAR mechanism remains untouched. Furthermore, we address a common scenario where the model is trained using multi-modal inputs but deployed on vehicles equipped with only one of the sensors. To assess robustness, we evaluate the model’s performance under conditions of complete sensor failure, where either the camera or LiDAR is missing.
@@ 3.4 BEV View Transformation
The body of works in BEV transformation can be divided based on the use of depth estimation techniques [6]. One faction, as discussed in [3, 22, 51, 65], embeds a distinct depth-estimation branch within their systems. Given the inherent challenges in predicting 3D bounding boxes from images, these models first forecast a per-pixel depth map. This map then serves as a compass, guiding image features to their rightful 3D coordinates. The subsequent BEV transformation process often follows a bottom-up approach, as depicted in [4]. On the other side of the spectrum are models leveraging pre-defined object queries [2, 4] or lean proposals [55, 54] to collate 2D features in a top-down manner. While both these paradigms have demonstrated their prowess on original datasets, we further expand the horizon by examining their efficacy on corrupted data.
@@ 4.1 Dataset Generation
We also consider scenarios simulating complete sensor blackouts in our camera-LiDAR fusion algorithms. Every pixel is set to zero when simulating the camera’s absence. To emulate the missing LiDAR readings, only the data points within a [[-[ -45, 45]]]] degree frontal field-of-view (FOV) are retained. Such a design choice is rooted in our observations that multi-modal trained models suffer from significant performance drops when LiDAR readings are entirely absent.
@@ 5.2 Camera-Only Benchmark
Our investigations further highlight a potential vulnerability in resilience rates across various corruptions. Even though the mCE metric displays a linear correlation between the nuScenes and nuScenes-C datasets, the mRR metric elucidates notable disparities among models with comparable baseline performance as shown in Table VII. This suggests potential overfitting of some models to the nuScenes dataset, thereby compromising their adaptability to the nuScenes-C dataset. For instance, despite Sparse4D [55] outperforming DETR3D [4] on the “clean” dataset, it falls short in terms of mRR metrics across all corruption categories. Moreover, DETR3D performs exceptionally well in Dark conditions, in stark contrast to BEVerse (swin-t) [51]. Despite its strong performance in clean conditions, BEVerse (swin-t) achieves only 12% relative performance in the dark. Therefore, a comprehensive assessment of cutting-edge models is essential for a complete evaluation of their capabilities.
As shown in Table II, to gain deeper insights into the model’s robustness, we dissected BEV models based on components like training strategies (e.g., pre-training and CBGS [46] resampling), model architectures (e.g., backbone), and learning techniques (e.g., temporal cue learning) and discuss them consequentially in Section 6.
@@
TABLE VI: Benchmark results for BEV detection robustness under sensor failures. The models are trained using multi-modal (camera and LiDAR) inputs while tested using single sensor inputs. For LiDAR failure, we keep the partial field of view in front of the ego vehicle since none of the candidate models can even operate with a complete LiDAR modality missing.
@@ 5.3.1 Camera Sensor Corruptions
We studied scenarios where cameras are impaired while LiDAR operates optimally, a frequent occurrence in real-world conditions. For instance, LiDAR point cloud capture remains largely unhampered by lighting variations, whereas camera captures can degrade under limited light. Intentionally, we excluded conditions like Snow and Fog, as they could introduce noise to both camera and LiDAR readings. Results of these studies are depicted in Table V. Interestingly, multi-modal fusion models maintain high performance even when the camera data is compromised. When provided with original LiDAR and degraded camera inputs, BEVFusion [66] consistently outperforms its LiDAR-only counterpart, with a notably higher NDS score of 0.6928, across most types of camera corruptions, except Dark. This affirms the efficacy of using LiDAR data even when the camera data is suboptimal.
@@ 5.4 Validity Assessment
Since the corruption images are synthesized digitally, it is important to study how closely they are compared to real-world corruption. To study the validity of synthesized images, we conducted two experiments, including the pixel distribution study and corruption-augmented training.
@@ 5.4.1 Pixel Distributions
Assuming that a corruption simulation is realistic enough to reflect real-world situations, the distribution of a corrupted “clean” set should be similar to that of the real-world corruption set. We validate this using ACDC [110], nuScenes [1], Cityscapes [112], and Foggy-Cityscapes [111], since these datasets contain real-world corruption data and clean data collected by the same sensor types from the same physical locations. We simulate corruptions using “clean” images and compare the distribution patterns with their corresponding real-world corrupted data. We do this to ensure that there is no extra distribution shift from aspects like sensor difference (e.g., FOVs and resolutions) and location discrepancy (e.g., environmental and semantic changes). The pixel histograms are computed jointly on all RGB channels. As illustrated in the Appendix, the pixel distributions of our synthesized images exhibit similar trends to those of real-world data, thereby showing the potential to use synthesized data to mimic real-world corruptions from a pixel qualitative analysis perspective.
@@ 5.4.2 Corruption-Augmented Training
Furthermore, if the corruption simulation is realistic enough to reflect real-world situations, a corruption-augmented model should achieve better generalizability than the “clean” model when tested on real-world corruption datasets. Also, the corruption-augmented model should also show better performance on the clean dataset. We validate this using nuScenes, nuScenes-Night, and nuScenes-Rain. We adopt FCOS3D [9] as the baseline and train the model with corruption augmentation. For nuScenes-Night and nuScenes-Rain, we train the model on the Day-train and Dry-train splits and evaluate on the Day-val, Night-val, Dry-val, and Rain-val splits. The results can be seen in Table VIII. We observe that using synthesized images as the data augmentation strategy successfully improves the cross-domain robustness. Specifically, in day-to-night domain transitions, we observe a significant performance drop from 0.3867 to 0.0854 in the baseline model due to the large domain gap. However, when trained with corruption augmentation, the model’s cross-domain performance improves by 45.8%, thereby validating the validity of our synthesized images.
@@ 5.5 Robustness Improvement
Building upon our experiment, we further study how to fundamentally improve the robustness of BEV perception models. In Subsection 5.5.1, we simply use the corruption as data augmentation as in Subsection 5.4.2. In Subsection 5.5.2, we resort to advanced CLIP [113] as the model backbone since CLIP is well-known to show promising out-of-distribution generality [116, 115].
@@
5.5.1 Limitation of Corruption Augmentations
We start with how to improve the robustness from a data-centric perspective. Specifically, we study if treating synthesized corruption data as a data augmentation alone is enough to solve the problem. To systematically assess the effectiveness of this approach, we apply corruption augmentation to five models within our benchmark, with results detailed in Table IX. Our findings indicate that corruption augmentation substantially enhances performance against semantic corruptions within our proposed dataset, particularly for corruption types that originally posed significant challenges to the models (e.g., Motion Blur, Snow). However, augmenting training data with sensor corruption scenarios, such as the omission of camera information, does not significantly improve robustness. This outcome underscores the limitation of improving robustness solely from a data perspective and highlights the necessity for future research to develop more sophisticated modules capable of handling incomplete input scenarios. However, it should be noted that when using our benchmark as a robustness assessment, synthesized data augmentation should not be a proper practice as it defeats the purpose of the benchmark to analyze robustness from a model perspective.
@@ 5.5.2 Foundation Model Backbone
We further analyze how to improve robustness from both data and model perspectives. Recent studies [116, 115] have demonstrated that foundation models, trained on internet-scale datasets in an unsupervised manner, exhibit notable generality compared to models trained on conventional datasets (e.g., ImageNet [117]). Inspired by these findings, we explore the potential to transfer the generality of foundation models to BEV perception tasks. Specifically, we investigate three distinct approaches for utilizing the CLIP backbone in our models: (1) freezing the backbone while training only the detector head; (2) fine-tuning both the backbone and the detector head; (3) freezing the backbone and training the detection head, then fine-tuning the whole model together, since previous work [118] shows that end-to-end fine-tuning, even though improve in-distribution performance, can compromise robustness on out-of-distribution dataset.
The three methods are shown in Figure 4 and the results are shown in Table  X. The first observation is that CLIP is not well optimized for BEV perception tasks, as indicated by the low clean performance when we freeze the CLIP backbone. Additionally, end-to-end fine-tuning with randomly initialized detection heads leads to minimal improvement. Interestingly, when applying corruption augmentation, the CLIP backbone shows little improvement to the baseline model. The improvement in mRR is only 0.56 and, when equipped with corruption augmentation, the mRR is even lower. Finally, our two-stage training can effectively improve the performance while also transferring the robustness of CLIP towards BEV perception tasks, especially with equipped with corruption augmentation. The robustness improvement surpasses the end-to-end fine-tuned CLIP by a notable margin. For instance, the NDS score under Dark, Fog, and Snow improve by 23.1%, 11.8%, and 15.8% respectively. However, even if we show it’s possible to improve robustness by leveraging both data and model, the purpose of RoboBEV is mainly to focus on model aspects instead of data. Therefore, it’s not a proper practice for future works to evaluate the RoboBEV benchmark with those synthesized corruption as augmentation.
@@ 6.2 Model Pre-Training
- Pre-training improves robustness across a wide range of semantic corruptions while does not help with temporal corruptions. The effectiveness of these strategies for improving model robustness is illustrated in Figure 6a and Figure 6b, where models that utilize pre-training largely outperform those not. For controlled comparison, we re-implement the BEVDet (r101) [3] model using the FCOS3D [9] checkpoint as initialization. Our results, presented in Figure 10, show that pre-training can significantly improve mRR across a wide range of corruptions (except Fog) even if it has lower “clean” NDS (0.3780 vs. 0.3877). Specifically, under Color Quant, Motion Blur, and Dark corruptions, the mRR metric improves by 22.5%, 17.2%, and 27.8%, respectively. It is worth noting that pre-training mainly improves most semantic corruptions and does not improve temporal corruptions. Even though, the pre-trained BEVDet still largely lags behind those depth-free counterparts. Therefore, we can conclude that pre-training together with the depth-free bird’s eye view transformation provides models with strong robustness. More recently, M-BEV [114] proposed a masked pre-training task to enhance robustness under incomplete sensor input. We compare a masked pre-trained PETR with the corresponding baseline. The results can be seen in Table XI. We find M-BEV can most effectively improve robustness towards incomplete sensor output, and reveal the promise of masked image modeling pre-train in BEV perception tasks.
@@ 6.3 Temporal Fusion
- Temporal fusion has the potential to yield better absolute performance under corruptions. Fusing longer temporal information largely helps with robustness. We are particularly interested in examining how models utilizing temporal information perform under temporal corruptions. We find SOLOFusion [65] which fuses wider and richer temporal information performs extremely well compared to its short-only and long-only versions. In terms of Camera Crash, the short-only and long-only versions have close resilience rate performance shown in Table VII (65.04 vs. 65.13). However, the fusion version improves to 70.73, which is the highest among all the candidate models. Similarly, the fusion version improves the resilience rate by almost 10% compared to the other two versions under Frame Lost corruption. Moreover, the RR metric of its long-only version outperforms its short-only counterpart on a wide range of corruption types, which indicates the great potential of utilizing longer temporal information.